A 10% alcohol solution is to be combined with a 40% alcohol solution to make 20 liters of a 25% alcohol solution. How many liters of each are needed?
Answer by lwsshak3(11628)

A 10% alcohol solution is to be combined with a 40% alcohol solution to make 20 liters of a 25% alcohol solution. How many liters of each are needed?
**
let x=amt of 10% solution needed
20-x=amt of 40% solution needed
..
10%x+40%(20-x)=25%*20
.10x+8-.4x=5
-.30x=-3
x=3/.3=10
20-x=10
amt of 10% solution needed=10 liters
amt of 40% solution needed=10 liters